OpenStudy (anonymous):

HELP ME OMG a particle travels along a straight line with its acceleration at time t seconds equal to (3t+2) m/s2. the particle has an initial positive velocity and travels 30m in the fourth second. Find the velocity of the object when t=5

OpenStudy (anonymous):

anyone?

hartnn (hartnn):

insufficient data

OpenStudy (ghazi):

integrate that function you'll get velocity at t=5...since velocity \[v=\int\limits_{0}^{t} adt\]

OpenStudy (anonymous):

insufficient data? whaaat. you're just giving up

OpenStudy (anonymous):

@ghazi i cant just integrate blindly, it will give me an unknown constant.

OpenStudy (anonymous):

@ghazi, doesnt that give you velocity = 85 m/s?

OpenStudy (anonymous):

@ mathsloser56, nah it wont, its a definite integral

OpenStudy (ghazi):

see limit of integration will be 0 to 5 and you'll get v= 47.5m/s at t=5

OpenStudy (ghazi):

after integrating, you'll get (3t^2/2)+2t....now put t=5

hartnn (hartnn):

ok....velocity=integral of acceleration....which gives,v=3t*t/2+2t+c...where c is constant...so they should have given conditions to find initial vel=c....but they gave values of displacement,for which we need to integrate once more and now there will be two unknowns...so two conditions required...r u following??

OpenStudy (anonymous):

@ghazi but thats assuming you start with 0 velocity. the question says it has positive velocity

OpenStudy (anonymous):

@hartnn yes. i guess so

OpenStudy (anonymous):

this is not yet solved.

OpenStudy (anonymous):

im tempted to spam this box to get the attention

OpenStudy (anonymous):

ok got the answer 52. turns out im the best around here after all ;p

OpenStudy (anonymous):

wouldnt you agree?

OpenStudy (anonymous):

\[\sqrt{-1} + 2^{3} + \sum_{}^{} + \pi \] and it was delicious

hartnn (hartnn):

i ate some pie :)